Q#5 An electron in a circular orbit about a proton can be described by classical mechanics if its angular momentum L is very much greater than h. Show that this condition is satisfied if the radius of the orbit r is very much greater than the Bohr radius a0 , i.e if

r>>a0 = 4πϵh2/e2me
1
Expert's answer
2020-06-22T11:10:45-0400

According to Bohr, the quantization of the angular momentum of electrons is


"L=mvr=n\hbar."

On the other hand, we know that the angular momentum, according to quantum mechanics, is


"L=\sqrt{l(l+1)}\hbar=\sqrt{n^2-n}\cdot\hbar."

Of course, as we see, the two equation are different:


"n>\sqrt{n^2-n}."

However, we can say that "n\approx\sqrt{n^2-n}" for very large values of "n". This criterion leads to "L>>\hbar". Therefore, according to the first equation ("L=mvr=n\hbar"), we see that this is only possible for very large values of "r".


Hence, to describe the motion of an electron in terms of classical mechanics, we need "r>>a_0". As we saw above, this is only possible for very big "n", which leads to "L>>\hbar".